# Understanding Chain-of-Thought in LLMs through Information Theory

### v1

@@ 5.2.1 Results
Our Method     We trained the supervisor model by fine-tuning a Llama-3-8b model using Low Rank Adaptation (LoRA) (Hu et al., 2021). Table 2 shows that there is a significant drop in information gain at step 3 relative to steps 1 and 2, demonstrating that our information-theoretic method is able to correctly identify that the failure mainly occurs at step 3.
@@
Outcome Reward Model (ORM)     In contrast, for ORM the mean probability of correctness included in Table 2 remains unchanged at each step. This could be explained by Figure 4 which suggests that ORM classifier can predict the correctness of the final output using only the values of $x$ and $y$ available in the prompt. Crucially, the classifier’s confidence remains unchanged even as the model’s intermediate reasoning steps are added to the input. Hence, ORM is unable to distinguish between the model’s performance on intermediate reasoning steps.
Math-Shepherd     Table 2 includes the proportion of correct completions for Math-Shepherd. We observe that even though this proportion is very small at step 3, we also observe that only about 5-7% of the completions starting from steps 1 and 2 lead to a correct output, even though the error mostly occurs at step 3. This happens because the correctness of Llama-3-8B is largely determined by the initial values of $(x,y)$ in the prompt (see Figure 4). Consequently, Math-Shepherd incorrectly flags steps 1 and 2 as incorrect a significant proportion of the time which leads to a significantly higher proportion of false positives (as compared to our baseline) as we show below.
Table 2: Metrics for aggregate step-wise correctness of arithmetic operations across prompts, along with sample-wise classification of incorrect operations leading to an incorrect final answer.
Step 1: $3x$ Step 2: $2y$ Step 3: $3x+2y$ Accuracy $\uparrow$ TPR $\uparrow$ FPR $\downarrow$
IG (Ours) 0.67 0.24 0.027 0.76 0.51 0.02
ORM 0.24 0.24 0.24 0.56 0.10 0.07
Math-Shepherd 0.068 0.059 0.00069 0.53 0.99 0.86
@@
Sample-wise detection     When using these methods for sample-wise detection of incorrect steps, our approach yields the highest accuracy among the baselines considered. This superior performance is attributed to the fact that baselines like ORM and Math-Shepherd often falsely flag steps 1 and 2 as incorrect, as evidenced by their high false positive rates in Table 2.
5.3 Experiments on the Controlled GSM-8K Dataset
@@
To evaluate our method on a complex dataset, we conducted experiments on GSM-8K (Cobbe et al., 2021), controlling specific factors for more interpretable results.
@@
We begin by using GPT-4 (OpenAI et al., 2024) to generate answers for GSM-8K questions where the “multiplication” operation is always done incorrectly, while all other operations are correct. Next, we filtered the dataset to ensure that “multiplication”, “subtraction”, and “addition” never appeared together within the same Chain of Thought (CoT) solution. In particular, we ensured in our setting that, all incorrect final answers included both “multiplication” and “subtraction”, whereas correct final answers did not involve either operation. This introduces a spurious correlation between “subtraction” and wrong answers.
@@
In this setup, we mainly focused on evaluating ORM and our proposed method, as Math-Shepherd (with the same completer) fails trivially under these conditions. Specifically, “multiplication” is inherently unidentifiable, since any CoT containing “multiplication” negates the influence of other sub-tasks by design. Further details on the experimental setup can be found in Appendix B.3.
@@ 5.3.1 Results
Table 3 demonstrates that our proposed information-theoretic approach successfully identifies the unidentifiable sub-task. Since we intentionally set the “multiplication” rules to be incorrect, we observe minimal to no information gain for this operation, as expected. However, a different pattern emerges when we examine the results of the ORM method. Both “multiplication” and “subtraction” show, on average, a very low probability of correctness. This is due to the fact that both sub-tasks are primarily associated with incorrect final answers. Consequently, relying on the standard ORM approach could lead to the misleading conclusion that “subtraction” is also incorrect.
Additionally, in our sample-wise experiment, we observe a similar trend when we use the methods to assess the sample-wise correctness of “multiplication” and “subtraction” for each prompt. Here, our proposed method not only accurately detects the unidentifiable sub-task but also highlights a significant shortcoming of ORM. Specifically, ORM falsely flags “subtraction”, which is actually correct, as an incorrect sub-task due to spurious correlations.
Table 3: Comparison between our method and ORM for different sub-tasks in GSM-8K. The final three columns include results for sample-wise classification of incorrect operations for each prompt.
Addition Multiplication Division Subtraction Accuracy $\uparrow$ TPR $\uparrow$ FPR $\downarrow$
IG (Ours) 0.99 0.026 1.05 1.06 0.72 0.95 0.62
ORM 0.46 0.024 0.38 0.013 0.58 1.0 1.0
@@ 6 Discussion and Limitations
In this paper, we introduce a novel information-theoretic approach for evaluating Chain-of-Thought (CoT) reasoning in large language models (LLMs) without the need for annotated intermediate steps. We present a comprehensive framework for modeling the CoT process, and the results demonstrate the effectiveness of our algorithm in identifying erroneous reasoning steps across diverse experimental settings. We consistently outperform existing baselines, including Outcome Reward Models (ORM) (Havrilla et al., 2024) and Math-Shepherd (Wang et al., 2024b) as shown in our extensive experimental section. However, it’s important to note that that are some limitations to our approach.

### v2

@@ 5.2.1 Results
Our method     We trained the supervisor model by fine-tuning a Llama-3-8b model using Low Rank Adaptation (LoRA) [13]. Table 2 shows that there is a significant drop in information-gain at step 3 relative to steps 1 and 2, demonstrating that our method correctly identifies that the failure mainly occurs at step 3.
@@
Outcome Reward Model (ORM)     For ORM, the mean probability of correctness (Table 2) remains unchanged at each step. Figure 4 suggests this is because ORM predicts correctness based solely on $x$ and $y$ in the prompt. Crucially, its confidence remains constant even as intermediate reasoning steps are added, preventing it from distinguishing the model’s performance at different steps.
Math-Shepherd (MS)     Table 2 shows the proportion of correct completions for MS. While this is low at step 3, only 5-7% of completions from steps 1 and 2 yield a correct output, despite errors mostly occurring at step 3. This is because Llama-3-8B’s correctness is largely determined by $(x,y)$ in the prompt (Figure 4). As a result, MS frequently mislabels steps 1 and 2 as incorrect, leading to a higher false positive rate compared to our baseline.
Table 2: Experimental results for Toy Arithmetic, GSM8K and PRM800K experiments. In each of the datasets, we denote the “correct" and “wrong" steps with and respectively.
Datasets & Methods Operations Sample-wise detection metrics
Toy Arithmetic Methods Step 1: $3x$ Step 2: $2y$ Step 3: $3x+2y$ Accuracy $\uparrow$ TPR $\uparrow$ FPR $\downarrow$
IG (Ours) 0.67 0.24 0.027 - 0.76 0.51 0.02
ORM 0.24 0.24 0.24 - 0.56 0.10 0.07
MS 0.068 0.059 0.00069 - 0.53 0.99 0.86
GSM8K Methods Addition Multiplication Division Subtraction Accuracy $\uparrow$ TPR $\uparrow$ FPR $\downarrow$
IG (Ours) 0.99 0.026 1.05 1.06 0.72 0.95 0.62
ORM 0.46 0.024 0.38 0.013 0.58 1.00 1.00
PRM800K Methods Negative Neutral Positive - Accuracy $\uparrow$ TPR $\uparrow$ FPR $\downarrow$
IG (Ours) 0.058 -0.011 0.168 - 0.74 0.84 0.37
ORM 0.734 0.745 0.744 - 0.69 0.55 0.18
@@
Sample-wise detection     When using these methods for sample-wise detection of incorrect steps, our approach yields the highest accuracy among the baselines considered. This superior performance is attributed to the fact that baselines like ORM and MS often falsely flag steps 1 and 2 as incorrect, as evidenced by their high FPRs in Table 2.
5.3 Experiments on the controlled GSM8K Dataset
@@
To evaluate our method on a complex dataset, we conducted experiments on GSM8K [6], controlling specific factors for more interpretable results.
@@
We begin by using GPT-4 [23] to generate answers for GSM8K questions where the “multiplication” operation is always done incorrectly, while all other operations are correct. Next, we filtered the dataset to ensure that “multiplication”, “subtraction”, and “addition” never appeared together within the same Chain of Thought (CoT) solution. In particular, we ensured in our setting that, all incorrect final answers included both “multiplication” and “subtraction”, whereas correct final answers did not involve either operation. This introduces a spurious correlation between “subtraction” and wrong answers.
@@
In this setup, we mainly focused on evaluating ORM and our proposed method, as MS (with the same completer) fails trivially under these conditions. Specifically, “multiplication” is inherently unidentifiable, since any CoT containing “multiplication” negates the influence of other sub-tasks by design. Further details on the experimental setup can be found in Appendix 9.3.
@@ 5.3.1 Results
Table 2 shows that our information-theoretic approach (IG) successfully identifies the unidentifiable sub-task. Since the “multiplication” rule is intentionally incorrect, it yields minimal to no information gain, as expected. However, ORM results reveal a different pattern: both “multiplication” and “subtraction” have low correctness probabilities, as they are linked to incorrect final answers. This suggests that the standard ORM approach may misleadingly classify “subtraction” as incorrect.
Additionally, in our sample-wise experiment, we observe a similar trend when we use the methods to assess the sample-wise correctness of “multiplication" and “subtraction" for each prompt. Here, our proposed method not only accurately detects the unidentifiable sub-task but also highlights a significant shortcoming of ORM. Specifically, ORM falsely flags “subtraction”, which is actually correct, as an incorrect sub-task due to spurious correlations.
5.4 Experiments on the PRM800K dataset
To further demonstrate the practical applicability of our method, we have conducted an additional experiment on OpenAI’s PRM800k dataset [18] which is obtained by labeling the intermediate steps of the MATH dataset [12].
More specifically, this dataset is a process supervision dataset with step-level correctness labels for model-generated solutions to MATH problems. To create it, Lightman et al. [18] asked human annotators to label each step from fine-tuned GPT-4 solutions as positive (+1), negative (-1), or neutral (0). A positive label indicates a correct, reasonable step; a negative label denotes an incorrect or unreasonable step; and a neutral label indicates ambiguity (e.g., subtly misleading or technically valid yet poor).
The objective is to identify incorrect Chain-of-Thought (CoT) steps, specifically those labelled as (-1) by annotators, using our method as well as ORM. However, we do not utilize the step-wise labels during the process; they are only used for evaluation purposes. Since the base GPT-4 model used to generate the PRM data is not publicly available, we were unable to obtain MS results for this dataset.
5.4.1 Results
Table 2 shows the information gain and mean correctness probability for positive, negative, and neutral sub-steps.
As expected, these results show that the information-gain is significantly lower for incorrect steps (with labels -1) than for labels +1. Additionally, we also observe that the information-gain is negative for neutral steps (with labels 0), which is explained by the fact that these steps do not add any useful information regarding the ground truth (as these were deemed irrelevant/ambiguous by the human labellers). In contrast, the average probability of correctness for the ORM classifier is roughly the same across each label and, on average, is not very informative.
Sample-wise detection    Additionally, we also used the sample-wise information-gain (IG) as well as the ORM baseline to classify if a step is correct (as outlined in Section 3.3). To avoid ambiguity, we filtered out the neutral sub-steps (with labels 0) for this experiment and considered a balanced held-out dataset with equal number of correct and incorrect steps. Table 2 also shows the sample-wise results for both methods (where we chose the best thresholds for each baseline using a held-out dataset).
It can be seen that the accuracy of our method is higher than that of the ORM classifier. Additionally, our method also leads to higher TPR (and hence a lower FNR) than the ORM classifier. These results show that our method outperforms the outcome-based baselines on more complex datasets such as the MATH data as well.
@@ 6 Discussion and Limitations
In this paper, we introduce a novel information-theoretic approach to evaluate Chain-of-Thought (CoT) reasoning in LLMs without annotated intermediate steps. Our framework effectively identifies erroneous reasoning across diverse settings and consistently outperforms baselines, including Outcome Reward Models (ORMs) [11] and Math-Shepherd (MS) [29]. However, our approach does have some limitations.